find the smallest positive solution of the system of linear congruences ? X=2 (mod3) , X=3(mod 5), X=2(mod 7)
1
Expert's answer
2020-10-21T15:57:32-0400

Since the numbers 3 5 7 are coprime, we use the Chinese remainder theorem

"{M_0} = {m_1}{m_2}{m_3}=3*5*7 =105"

"M_i =\frac{M_0}{m_i}"

"M_1 =\frac{M_0}{m_1}=\frac{105}{3}=35"

"M_2 =\frac{M_0}{m_2}=\frac{105}{5}=21"

"M_3 =\frac{M_0}{m_3}=\frac{105}{7}=15"

"M_iy_i =a_i(mod\ m_i)"


"35y_1 =2(mod\ 3)"

"33y_1+2y_1 =2(mod\ 3)"

"2y_1 =2(mod\ 3)"

"y_1=1"


"21y_2 =3(mod\ 5)"

"20y_2+y_2 =3(mod\ 5)"

"y_2 =3(mod\ 5)"

"y_2 =8"


"15y_3=2(mod\ 7)"

"14y_3+y_3=2(mod\ 7)"

"y_3=2(mod\ 7)"

"y_3=9"

"x =( M_1y_1+M_2y_2+M_3y_3)(mod\ M_0)"

"x =(35*1+21*8+15*9)(mod\ 105)"

"x= 338(mod\ 105)"

"x=23"

Answer: x=23